OpenStudy (atomicreaper):

If AC = 5 cm, BC = 12 cm, and , the radius of the circumscribed circle is cm and mOAC°.

OpenStudy (xmissalycatx):

Send a picture?

OpenStudy (xmissalycatx):

Okay, well to find the length of AB just use pythagorean theorem which is your diameter, then cut that in half..

OpenStudy (atomicreaper):

So how would I set that up? I know its a^2+b^2=c^2, but how would I set it up since its a circle

OpenStudy (xmissalycatx):

You would just use the measurements of the triangle..

OpenStudy (xmissalycatx):

so 5^2 + 12^2 = c^2

OpenStudy (xmissalycatx):

And then divide the square root of your c by 2 because you need the radius not diameter.

OpenStudy (atomicreaper):

so it would be 6.5?

OpenStudy (xmissalycatx):

I didn't solve for it, but that seems to be about right. :P

OpenStudy (atomicreaper):

Ok so what about the second part, what do I do for it?

OpenStudy (xmissalycatx):

http://www.teachertube.com/video/solve-for-angle-in-triangle-in-circle-80962 watch this. I'm pretty sure angle C is 90 degrees btw.

OpenStudy (atomicreaper):

Thank you so much !!!

OpenStudy (xmissalycatx):

You're very welcome! :D